OpenStudy (anonymous):

Find the gradient of the function and the maximum value of the directional derivative at the given point. h(x,y)=ycos(x-y) (0,pi/3)

OpenStudy (anonymous):

I want to see if I've done this right. So taking the partial derivative, I get \[\Delta h(x,y)=-ysin(x-y)i+\cos(x-y)+ysin(x-y)\]

OpenStudy (anonymous):

and then evaluating at (0,pi/3) would get me:\[ (-\pi/3)(-\sqrt{3}/2)+(-\sqrt{3}/2)+(\pi/3)(-\sqrt{3}/2)\]

OpenStudy (anonymous):

Gradient meaning U've got to fing dy/dx is it????????

OpenStudy (anonymous):

Yeah, you have to take the partial derivative of each term: \[\Delta f(x,y)=f_{x}(x,y)i+f _{y}(x+y)j\]

OpenStudy (anonymous):

gfradient is (fx (x,y),fy(x+y)) is a vector. If they asking you for maximum value, i would say they asking either for modul or just the value of partial derivatives at the point